OpenStudy (anonymous):

What is 1000000100100110 in decimal, hexadecimal, and base 36?

OpenStudy (kropot72):

The value in decimal is found from the following: \[2^{1}+2^{2}+2^{5}+2^{8}+2^{15}=\] Can you calculate the answer?

OpenStudy (kropot72):

@aetteews Are you there?

OpenStudy (anonymous):

Yup, sorry, was taking awhile for people to respond... 2+2^2+2^3+2^5+2^8+2^15 = 33062

OpenStudy (kropot72):

You are correct

OpenStudy (anonymous):

Okay, so that gets me to decimal...

OpenStudy (anonymous):

So how do I convert to hexadecimal?

OpenStudy (anonymous):

Okay... so 16^3 = 4096 and 16^4 = 65536... so 16^3 is less than 33062. So then 33062 - 16^3 = 28966.

OpenStudy (anonymous):

Okay, I think I remember this, do you use the modulus (%) operator for this part?

OpenStudy (kropot72):

First find the highest power of 16 that will divide into 33062. Then find the highest number that can be multiplied with 4096 to be less than 33062 (ignore the second step I gave before)

OpenStudy (anonymous):

Divide into? You mean evenly?

OpenStudy (anonymous):

So 33062/16^3= 8.07177734

OpenStudy (kropot72):

No. Not evenly. In this case there will be a remainder that will be treated next.

OpenStudy (anonymous):

Okay, so the remainder can be calculuted using modulus. So the remainder is 294.

OpenStudy (kropot72):

So 8 is the most significant digit of the hexadecimal answer. 16^2 will divide into 294 with remainder 38. So 1 is the next digit of our solution.

OpenStudy (anonymous):

Oh, okay, so then you do the same thing again? 294/16^2=1.1484375 294%16^2= 38 38/16=2.37500 38%16=6 What now?

OpenStudy (kropot72):

The full hex number is 8126. Can you see how it was obtained.

OpenStudy (anonymous):

Yup, it's the 8 from the first part 33062/16^3= 8.07177734 the 1 is from 294/16^2=1.1484375 and the 2 is from 38/16=2.37500, so I guess the remainder once you can't divide out anymore becomes your last digit?

OpenStudy (kropot72):

You are right. The least significant digit is one of the 16 hex numbers 0,1, 2, 3, 4, 5, 6, 7, 8, 9, A, B, C, D, E, F. The least significant digit is simply the number of units (16^0 = 1). Sorry I need to go now. Perhaps someone else can help you now.

OpenStudy (anonymous):

I think I've got it, thanks for the help! :)

OpenStudy (kropot72):

Great. I enjoyed working with you :)